OpenStudy (moonlitfate):

A car traveling at 32 m/s passes a policeman wo immediately accelerates at 2.43 m/s^2 from rest in pursuit. To the nearest second how long does it take the policeman to catch the speeder?

OpenStudy (abhisar):

Well do you know the answer of this question ?

OpenStudy (moonlitfate):

I don't, and I'm trying to see if any of the formulas that you gave would work here.

OpenStudy (abhisar):

Well it's a tricky one ! You'll have to use one of the formulas

OpenStudy (moonlitfate):

Well, we have acceleration (a) for two cars...

OpenStudy (abhisar):

Just imagine the situation, as soon as the policemen sees the car, he starts following it and finally catches it. Here, distance covered and time taken by the car and the policemen during the pursuit will be same

OpenStudy (abhisar):

The acceleration given is just for the policemen's car [if he is chasing by a car :p]. The other car is moving with a constant velocity of 32 m/s

OpenStudy (moonlitfate):

Confusing, lol.

OpenStudy (moonlitfate):

Hmm, would we have to use the first one?

OpenStudy (moonlitfate):

v = u + at ?

OpenStudy (moonlitfate):

And solve ti for t?

OpenStudy (abhisar):

Let t be the time taken which would be same for both. Now, \(\sf 32 \times t=0\times t+2.43\times t^2\) Now solve it for t

OpenStudy (abhisar):

?

OpenStudy (moonlitfate):

Still trying to figure it out. It's onl proving that I need to review basic Algebra...

OpenStudy (abhisar):

Let t be the time, then distance traveled by first car = 32*t [in uniform motion distance=speed*time] distance traveled by police = 0*t+1/2*2.43*\(\sf t^2\) Since distance covered during the pursuit by both the cars would be same we can equate the two equations and write it as \(\sf 32 \times t=0\times t+2.43\times t^2\) Solving it for t we get 32/2.43=t t=13.16 s

OpenStudy (abhisar):

is it clear @MoonlitFate ?

OpenStudy (moonlitfate):

Yeah, I didn't notice that the equation was a quadratic. :( So, I was making things so much more difficult than they needed to be.